### v1

@@ 1.2 Previous Work
A versatile approach for a robust PD solution is arguably to operate at the block layer, whereby the scheme exposes a block device interface (a common abstraction layer used by many operating systems to represent storage devices as arrays of fixed-size data blocks), providing just a bRead and a bWrite method. This is the approach used by solutions like TrueCrypt, and also by Shufflecake; in the remainder of this work, we will only focus on block-layer solutions. In this framework, a single underlying disk or device is formatted to host one or more volumes (logical storage units, usually represented as virtual block devices), each encrypted with a different (usually password-derived) symmetric key. When confronted with the adversary, the user surrenders the passwords to some of these volumes, called decoy volumes (because in some use cases they might contain deceptively innocent data): PD security guarantees that, even after these passwords have been given up, inspection of the disk by the adversary still yields no clue whatsoever hinting at the presence of some further, hidden volumes. The intuition behind the formal definition of these guarantees, is that the adversary cannot distinguish between the case where all the passwords have been surrendered, and the case where there is still undisclosed secret information left.
@@
Multi-snapshot attacks are a well-known issue in PD systems (TrueCrypt and derivatives are also vulnerable in this sense) and designing countermeasures turns out to be challenging. As of today, only a few constructions [10] achieve multi-snapshot security, but at a hefty performance cost that makes them not practical for most use cases. Most of these solutions are based on oblivious random access machines (ORAMs). ORAMs are cryptographic schemes that aim at hiding the access patterns (in addition to the data content itself) of a trusted agent accessing an untrusted storage. The connection between ORAMs and PD has been investigated since 2014 with the HiVE construction [6]. In a nutshell, the idea is that if we use an ORAM to access a device, then nobody, not even a run-time backdoor in the device firmware, can know which (logical) location we access and how, thereby providing a solid method for implementing PD. However, ORAMs are extremely slow: it is known [26] that the bandwidth overhead of any secure ORAM of size $n$ is $\Omega(\log(n))$. The HiVE paper circumvented this problem with the following observation: If we are not worried by run-time backdoors in the device firmware, but are only concerned about “traditional” multi-snapshot adversaries, i.e. post-arrest investigation of the device physical layer, then we do not need a fully-fledged ORAM, because read operations do not change the state of the device. So all we need is a “write-only” ORAM (WoORAM) that only obfuscates write requests. The advantage is that there is no known efficiency bound for WoORAMs, and in fact existing WoORAM constructions seem to be slightly better than full ORAMs. The WoORAM approach sparked a whole new line of research in multi-snapshot resistant PD solutions [39, 9], and it has been proven [10] that resistance to multi-snapshot security under certain assumptions is equivalent to the use of WoORAMs.
@@ 1.3 Limitations of Existing Solutions
Moreover, WoORAMs solutions themselves might not be bulletproof. In fact, we believe that the idea that read requests do not change the underlying state of the physical device is a somewhat strong assumption, and hard to justify with modern, complex SSDs that might, for example, cache read requests in some undocumented memory area of the firmware, or store read data on an ad-hoc buffer to improve performance, etc.
@@ 2 Preliminaries
In this section we give the required preliminaries that are going to be used in the rest of this work. In the following, we use “iff” as “if and only if”. Array indices start from $1$. By (efficient) algorithm or procedure we mean a uniform family of circuits of depth and width polynomial in the index of the family. We implicitly assume that all algorithms take the index of the family as a first input, so we will often omit this. In the case of cryptographic algorithms, we call such index a security parameter, and we denote it by $\lambda$. We will often label algorithms with names that reflect their role, e.g. “adversary”, “distinguisher”, etc. If an algorithm $A$ is deterministic, we denote its output $y$ on input $x$ as $y:=A(x)$, while if it is randomized we use $y\leftarrow A(x)$; when derandomising an algorithm we look at the deterministic algorithm obtained when considering explicitly the internal randomness $r$ as an additional auxiliary input, and we write $y:=A(x;r)$ We will call negligible (and denote by $\mathsf{negl}(x)$) a function that grows more slowly than any inverse polynomial in $x$, and overwhelming a function which is $1$ minus a negligible function. Given an event $E$, we denote by $\bar{E}$ its negation. Finally, we will write $x\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}X$ if $x$ is sampled uniformly at random from a set $X$.
@@ 2.1 Cryptographic Primitives
[…]
@@
Regarding encryption, the most fundamental primitive is symmetric-key encryption (SKE), also called secret-key encryption. An SKE scheme is a pair of algorithms (one for encryption and one for decryption) which define a bijection between a domain (plaintext space) and a co-domain (a subset of the ciphertext space) of strings. An additional input, the secret key, fixes the bijection across the set of all possible ones, and correctness of the SKE ensures that, if the same secret key is used, then the bijection offered by the decryption algorithm is the inverse of that offered by encryption (with overwhelming probability in the case of non-deterministic decryption). The size of a typical secret key in many real-world applications is 128 or 256 bit. This allows to index at most $2^{128}$ or $2^{256}$ unique bijections, which is generally much smaller than the number of possible bijections as the domain space gets larger. For this reason, it is generally impossible to ask the coverage of all possible bijections as a security property of SKEs. Instead, security is usually given in terms of indistinguishability games, with strength of the resulting notions depending on the additional power granted to the adversary. One of the most common security notions for SKEs is indistinguishability under chosen plaintext attack (or $\mathsf{IND{\mbox{-}}CPA}$ in short). In such game, the adversary’s goal is to distinguish between the encryption of two messages of her choice, given additional access to an encryption oracle (for the same, unknown secret key used in the game). The scheme is called $\mathsf{IND{\mbox{-}}CPA}$ secure iff no efficient adversary $\mathcal{A}$ can successfully win with probability more than negligibly better than guessing at random. Notice that this, in particular, requires that the SKE must be randomised, i.e., encrypting twice the same plaintext with the same key will generally yield two different ciphertexts. We will denote encryption (resp., decryption) of a plaintext $p$ (resp., ciphertext $c$) with a key $k$ (and, optionally, a randomness $r$) as $\texttt{Encrypt}(p,k;r)$ (resp., $\texttt{Decrypt}(c,k;r)$). The randomness $r$ is generally not needed for decryption, so we will omit it in that case.
@@
As a building block for SKEs, block ciphers are widely used. These are algorithms that typically offer two different interfaces, one for encryption and one for decryption. In encryption mode, they take as input a block of plaintext of fixed size $B$ (the block size) and an encryption key of $K$ bits (the key size) and return a block of ciphertext, also of size $B$. This mapping is undone in decryption mode, provided the same key is used. In other words, block ciphers implement a subset of size at most $2^{K}$ of the space of all possible $(2^{B})!$ permutations (and their inverses) over $B$-bit strings. One of the most widely used block ciphers are those from the AES family [20], identified by AES-$K$ (with a block size of 128 bits, and a keysize $K$ of 128, 192, or 256 bits).
@@ 2.3 Plausible Deniability
In this section, we present a formal game-based definition of PD security. It is worth noting that almost every paper in the field has given its own security definition, always slightly different from the others; valid attempts have been made to unify them into a single framework [10], but here we will follow the arguably more intuitive one given in [6], which is well suited for the block-layer scenario we work in. In this setting, a user employs a PD scheme to multiplex a single storage device into $\ell$ independent volumes $V_{1},V_{2},\ldots V_{\ell}$, each $V_{i}$ being associated to a different password $P_{i}$. The PD scheme supports up to $\mathsf{max}$ volumes per device (so $1\leq\ell\leq\mathsf{max}$); the value of $\mathsf{max}$ is publicly known. Both the volumes and the underlying device are block-addressable, meaning that the read and write operations they support have the granularity of a block. The semantics of a scheme $\Pi$ is given as follows.
@@
Let $\ell\leq\mathsf{max}$, and $P_{1},\ldots P_{\ell}$ user-provided passwords. A PD scheme $\Pi$ is a tuple of algorithms:
  • $\Pi.\mathtt{setup}(P_{1},\ldots,P_{\ell})\to\Sigma$: Initialises the disk to host $\ell$ volumes $V_{1},\ldots V_{\ell}$, encrypted with passwords $P_{1},\ldots P_{\ell}$; returns a device instance description $\Sigma$ which encapsulates everything.
  • $\Pi.\mathtt{read}(\Sigma,i,B)\to d$: Reads data $d$ in block address $B$ from volume $V_{i}$ (we assume $\mathtt{read}$s to not modify the instance) [^2] Note this is also the case for WoORAM-based constructions, but not necessarily for ORAM-based or other, arbitrary, PD schemes constructions..
  • $\Pi.\mathtt{write}(\Sigma,i,B,d)\to\Sigma^{\prime}$: Writes data $d$ into block address $B$ of volume $V_{i}$, and updates the instance.
The following correctness requirement applies: for any fixed block $B$ and volume $V_{i}$, if $\,\Pi.\mathtt{write}(\Sigma,i,B,d)$ is the most recent write query which precedes a query $\Pi.\mathtt{read}(\Sigma^{\prime},i,B)$, then $\Pi.\mathtt{read}(\Sigma^{\prime},i,B)\to d$ (for simplicity, we consider operations atomic).
@@
Let us define an access as the tuple $o=(\mathtt{op},i,B,d)$, with $\mathtt{op}\in\set{\mathtt{read},\mathtt{write}}$ (if $\mathtt{op}=\mathtt{read}$, then $d$ is the return value), and $i$ being the index of the volume targeted by the access. Let us also define an access pattern as a (chronologically) ordered sequence of accesses $O=<o_{1},\ldots,o_{n}>$. An empty access $o=\,\perp$ is also defined, which is simply ignored by the instance $\Sigma$.
@@ 2.3.2 PD security
The security game for PD inherits some high-level concepts from the $\mathsf{IND{\mbox{-}}CPA}$ game (ciphertext indistinguishability under chosen-plaintext attack) for secret-key encryption. The adversary is a distinguisher, and is challenged with deducing whether she is interacting with a $\Sigma$ encapsulating $\ell$ or $\ell-1$ volumes. Also, she is allowed to choose the $\mathtt{read}$ or $\mathtt{write}$ operations to be executed, to capture the idea that indistinguishability must hold no matter the accesses performed on the volumes, hence including adversarial ones. A secret bit $b$ within the game determines whether $\Sigma_{0}$ (containing $\ell$ volumes) or $\Sigma_{1}$ (containing $\ell-1$ volumes) is first instantiated in the game and made to interact with the adversary. In both cases, we allow the adversary to choose the first $\ell-1$ passwords [^3] This represents the most unfavourable situation for the user, as we consider these passwords compromised anyway., and her goal is to guess $b$.
@@
    $\mathcal{A}$ chooses $\ell\leq\mathsf{max}$ and chooses $\ell-1$ passwords $P_{1},\ldots,P_{\ell-1}$.
@@
    A secret random bit $b\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\{0,1\}$ is drawn. If $b=0$, then an additional secret high entropy password is sampled $P_{\ell}\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\{0,1\}^{\lambda}$, where $\lambda$ is the security parameter [^4] We abuse notation by representing a password as a binary string, but w.l.o.g. it is equivalent to the case of a user-chosen password with $\lambda$ bits of entropy. We assume that the user will choose a high entropy password at least for the hidden volume..
@@
    The game creates $\ell-b$ volumes: $\Pi.\mathtt{setup}(P_{1},\ldots,P_{\ell-b})\to\Sigma_{b}$
@@
      $\mathcal{A}$ chooses access patterns $O_{0}$ and $O_{1}$, where $O_{0}$ is, in the adversary’s intentions, aimed at $\Sigma_{0}$ (thus potentially containing some operations on $V_{\ell}$) and $O_{1}$ is aimed at $\Sigma_{1}$ (and so only contains operations on $V_{1},\ldots,V_{\ell-1}$).
@@
These constraints define the adversary goal, by specifying which two exact situations she has to distinguish between: if a PD scheme is secure (i.e., the adversary cannot distinguish) under the game enforcing such a constraint, the implication is that a user, having performed some access pattern $O_{0}$ including some operations on $V_{\ell}$, can plausibly claim to instead have executed a corresponding $O_{1}$, which only accesses the volumes $V_{1},\ldots,V_{\ell-1}$ (whose passwords have already been surrendered).
@@
Let us first clarify that some constraint is necessary in order to have any hope in the PD game against the adversary. Otherwise, the adversary could submit $O_{0}$ and $O_{1}$ containing completely different (logical) $\mathtt{write}$ accesses to the decoy volumes $V_{1},\ldots,V_{\ell-1}$, and there would be no way of making the two outcomes indistinguishable, since the adversary holds the passwords to those volumes, so it could trivially verify which of the two patterns was executed. This suggests the need for a minimal rule, stating that, whether $O_{0}$ or $O_{1}$ is executed, the resulting logical contents of the decoy volumes $V_{1},\ldots,V_{\ell-1}$ must be the same. From the user’s perspective, this basic requirement means that we do not try to disguise the $\mathtt{write}$s to decoy volumes as something else, both because we do not need to, and because there would be no way of doing it even if we wanted to.
Furthermore, we notice that many PD solutions (including those based on WoORAMs) treat $\mathtt{write}$ requests in a completely different way than $\mathtt{read}$ requests, i.e. a $\mathtt{write}$ request could trigger allocation or reshuffling of a certain volume sector. This might happen without breaking the minimal rule described above: if an adversary first $\mathtt{read}$s at a previously unallocated position $B$, obtaining data $d$, and then $\mathtt{write}$s the same data $d$ at the same location $B$, this might cause a change of state in the instance without changing its logical content (as mandated by the minimal rule above). This, in turn, would enable an attack where the adversary merely checks whether this state change happened or not. However, we must consider this attack trivial: detecting this kind of state change on a decoy volume would actually not compromise security. Therefore, in order to capture this concept, we also demand that the set of all blocks that are touched by $\mathtt{write}$ requests be the same for both $O_{0}$ and $O_{1}$, for all volumes $V_{1},\ldots V_{\ell-1}$. We stress that this extra constraint is also very minimal and does not weaken the security guarantees offered in practice; in fact, most schemes demand even stricter constraints, for example by demanding that $O_{0}$ and $O_{1}$ are of the same length, which we do not require here.

### v2

@@ 1.2 Previous Work
A versatile approach for a robust PD solution is arguably to operate at the block layer, whereby the scheme exposes a block device interface (a common abstraction layer used by many operating systems to represent storage devices as arrays of fixed-size data blocks), providing just a $\mathtt{bRead}$ and a $\mathtt{bWrite}$ method. This is the approach used by solutions like TrueCrypt, and also by Shufflecake; in the remainder of this work, we will only focus on block-layer solutions. In this framework, a single underlying disk or device is formatted to host one or more volumes (logical storage units, usually represented as virtual block devices), each encrypted with a different (usually password-derived) symmetric key. When confronted with the adversary, the user surrenders the passwords to some of these volumes, called decoy volumes (because in some use cases they might contain deceptively innocent data): PD security guarantees that, even after these passwords have been given up, inspection of the disk by the adversary still yields no clue whatsoever hinting at the presence of some further, hidden volumes. The intuition behind the formal definition of these guarantees, is that the adversary cannot distinguish between the case where all the passwords have been surrendered, and the case where there is still undisclosed secret information left.
@@
Multi-snapshot attacks are a well-known issue in PD systems (TrueCrypt and derivatives are also vulnerable in this sense) and designing countermeasures turns out to be challenging. As of today, only a few constructions [10] achieve multi-snapshot security, but at a hefty performance cost that makes them not practical for most use cases. Most of these solutions are based on oblivious random access machines (ORAMs). ORAMs are cryptographic schemes that aim at hiding the access patterns (in addition to the data content itself) of a trusted agent accessing an untrusted storage. The connection between ORAMs and PD has been investigated since 2014 with the HiVE construction [6]. In a nutshell, the idea is that if we use an ORAM to access a device, then nobody, not even a run-time backdoor in the device firmware, can know which (logical) location we access and how, thereby providing a solid method for implementing PD. However, ORAMs are extremely slow: it is known [26] that the bandwidth overhead of any secure ORAM of size $n$ is $\Omega(\log(n))$. The HiVE paper circumvented this problem with the following observation: If we are not worried by run-time backdoors in the device firmware, but are only concerned about “traditional” multi-snapshot adversaries, i.e. post-arrest investigation of the device physical layer, then we do not need a fully-fledged ORAM, because $\mathtt{read}$ operations do not change the state of the device. So all we need is a “write-only” ORAM (WoORAM) that only obfuscates $\mathtt{write}$ requests. The advantage is that there is no known efficiency bound for WoORAMs, and in fact existing WoORAM constructions seem to be slightly better than full ORAMs. The WoORAM approach sparked a whole new line of research in multi-snapshot resistant PD solutions [39, 9], and it has been proven [10] that resistance to multi-snapshot security under certain assumptions is equivalent to the use of WoORAMs.
@@ 1.3 Limitations of Existing Solutions
Moreover, WoORAMs solutions themselves might not be bulletproof. In fact, we believe that the idea that $\mathtt{read}$ requests do not change the underlying state of the physical device is a somewhat strong assumption, and hard to justify with modern, complex SSDs that might, for example, cache $\mathtt{read}$ requests in some undocumented memory area of the firmware, or store read data on an ad-hoc buffer to improve performance, etc.
@@ 2 Preliminaries
In this section we give the required preliminaries that are going to be used in the rest of this work. In the following, we use “iff” as “if and only if”. Array and sequence indices start from 00. By (efficient) algorithm or procedure we mean a uniform family of circuits of depth and width polynomial in the index of the family. We implicitly assume that all algorithms take the index of the family as a first input, so we will often omit this. In the case of cryptographic algorithms, we call such index a security parameter, and we denote it by $\lambda$. We will often label algorithms with names that reflect their role, e.g. “adversary”, “distinguisher”, etc. If an algorithm $A$ is deterministic, we denote its output $y$ on input $x$ as $y:=A(x)$, while if it is randomized we use $y\leftarrow A(x)$; when derandomising an algorithm we look at the deterministic algorithm obtained when considering explicitly the internal randomness $r$ as additional auxiliary input, and we write $y:=A(x;r)$. We will call negligible (and denote by $\mathsf{negl}(x)$) a function that grows more slowly than any inverse polynomial in $x$, and overwhelming a function which is $1$ minus a negligible function. Given an event $E$, we denote by $\bar{E}$ its negation. Finally, we will write $x\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}X$ if $x$ is sampled uniformly at random from a set $X$.
@@ 2.1 Cryptographic Primitives
$\mathbf{Adv}_{\Pi}^{\mathsf{Game}}\left(\mathcal{A}\right):=\left|\underset{}{\mathsf{Pr}}\left[{\mathcal{A}\left(\Pi,\mathsf{Game}\right)\to\mathsf{win}}\right]-\underset{}{\mathsf{Pr}}\left[{\mathsf{Guess}\left(\Pi,\mathsf{Game}\right)\to\mathsf{win}}\right]\right|\leq\mathsf{negl}.$
@@
Regarding encryption, the most fundamental primitive is symmetric-key encryption (SKE), also called secret-key encryption. An SKE scheme is a pair of algorithms (one for encryption and one for decryption) which define a bijection between a domain (plaintext space) and a co-domain (a subset of the ciphertext space) of strings. An additional input, the secret key, fixes the bijection across the set of all possible ones, and correctness of the SKE ensures that, if the same secret key is used, then the bijection offered by the decryption algorithm is the inverse of that offered by encryption (with overwhelming probability in the case of non-deterministic decryption). The size of a typical secret key in many real-world applications is 128 or 256 bit. This allows to index at most $2^{128}$ or $2^{256}$ unique bijections, which is generally much smaller than the number of possible bijections as the domain space gets larger. For this reason, it is generally impossible to ask the coverage of all possible bijections as a security property of SKEs. Instead, security is usually given in terms of indistinguishability games, with strength of the resulting notions depending on the additional power granted to the adversary. One of the most common security notions for SKEs is indistinguishability under chosen plaintext attack (or $\mathsf{IND{\mbox{-}}CPA}$ in short). In such game, the adversary’s goal is to distinguish between the encryption of two messages of her choice, given additional access to an encryption oracle (for the same, unknown secret key used in the game). The scheme is called $\mathsf{IND{\mbox{-}}CPA}$ secure iff no efficient adversary $\mathcal{A}$ can successfully win with probability more than negligibly better than guessing at random. Notice that this, in particular, requires that the SKE must be randomised, i.e., encrypting twice the same plaintext with the same key will generally yield two different ciphertexts. We will denote encryption (resp., decryption) of a plaintext ${p}$ (resp., ciphertext ${c}$) with a key ${k}$ (and, optionally, a randomness ${r}$) as $\mathtt{Encrypt}({p},{k};{r})$ (resp., $\mathtt{Decrypt}({c},{k};{r})$). The randomness ${r}$ is generally not needed for decryption, so we will omit it in that case.
@@
As a building block for SKEs, block ciphers are widely used. These are algorithms that typically offer two different interfaces, one for encryption and one for decryption. In encryption mode, they take as input a block of plaintext of fixed bitsize $S_{\mathsf{B}}$ (the block size) and an encryption key of $K$ bits (the key size) and return a block of ciphertext, also of size $S_{\mathsf{B}}$. This mapping is undone in decryption mode, provided the same key is used. In other words, block ciphers implement a subset of size at most $2^{K}$ of the space of all possible $(2^{S}_{\mathsf{B}})!$ permutations (and their inverses) over $S_{\mathsf{B}}$-bit strings. One of the most widely used block ciphers are those from the AES family [20], identified by AES-$K$ (with a block size of 128 bits, and a keysize $K$ of 128, 192, or 256 bits).
@@ 2.3 Plausible Deniability
In this section, we present a formal game-based definition of PD security. It is worth noting that almost every paper in the field has given its own security definition, always slightly different from the others; valid attempts have been made to unify them into a single framework [10], but here we will follow the arguably more intuitive one given in [6], which is well suited for the block-layer scenario we work in. In this setting, a user employs a PD scheme to multiplex a single storage device into $N$ independent volumes $V_{0},V_{1},\ldots V_{N-1}$, each $V_{i}$ being associated to a different password $P_{i}$. The PD scheme supports up to $N_{\mathsf{MAX}}$ volumes per device (so $2\leq N\leq N_{\mathsf{MAX}}$); the value of $N_{\mathsf{MAX}}$ is publicly known. Both the volumes and the underlying device are block-addressable, meaning that the $\mathtt{read}$ and $\mathtt{write}$ operations they support have the granularity of a block. The semantics of a scheme $\Pi$ is given as follows.
@@
Let $N\leq N_{\mathsf{MAX}}$, and $P_{0},\ldots P_{N-1}$ user-provided passwords. A PD scheme $\Pi$ is a tuple of algorithms:
  • $\Pi.\mathtt{setup}(P_{0},\ldots,P_{N-1})\to\Sigma$: Initialises disk to host $N$ volumes $V_{0},\ldots V_{N-1}$, encrypted with passwords $P_{0},\ldots P_{N-1}$; returns a device instance description $\Sigma$ which encapsulates everything.
  • $\Pi.\mathtt{read}(\Sigma,i,\beta_{\mathsf{L}})\to d$: Reads data $d$ in block address $\beta_{\mathsf{L}}$ from volume $V_{i}$ (we assume $\mathtt{read}$s to not modify the instance) [^2] Note this is also the case for WoORAM-based constructions, but not necessarily for ORAM-based or other, arbitrary, PD schemes constructions..
  • $\Pi.\mathtt{write}(\Sigma,i,\beta_{\mathsf{L}},d)\to\Sigma^{\prime}$: Writes data $d$ into block address $\beta_{\mathsf{L}}$ of volume $V_{i}$, and updates the instance.
The following correctness requirement applies: for any fixed block $\beta_{\mathsf{L}}$ and volume $V_{i}$, if $\,\Pi.\mathtt{write}(\Sigma,i,\beta_{\mathsf{L}},d)$ is the most recent write query which precedes a query $\Pi.\mathtt{read}(\Sigma^{\prime},i,\beta_{\mathsf{L}})$, then $\Pi.\mathtt{read}(\Sigma^{\prime},i,\beta_{\mathsf{L}})\to d$ (for simplicity, we consider operations atomic).
@@
Let us define an access as the tuple $o=(\mathtt{op},i,\beta_{\mathsf{L}},d)$, with $\mathtt{op}\in\set{\mathtt{read},\mathtt{write}}$ (if $\mathtt{op}=\mathtt{read}$, then $d$ is the return value), and $i$ being the index of the volume targeted by the access. Let us also define an access pattern as a (chronologically) ordered sequence of accesses $O=<o_{0},\ldots,o_{n-1}>$. An empty access $o=\,\perp$ is also defined, which is simply ignored by the instance $\Sigma$.
@@ 2.3.2 PD security
The security game for PD inherits some high-level concepts from the $\mathsf{IND{\mbox{-}}CPA}$ game (ciphertext indistinguishability under chosen-plaintext attack) for secret-key encryption. The adversary is a distinguisher, and is challenged with deducing whether she is interacting with a $\Sigma$ encapsulating $N$ or $N-1$ volumes. Also, she is allowed to choose the $\mathtt{read}$ or $\mathtt{write}$ operations to be executed, to capture the idea that indistinguishability must hold no matter the accesses performed on the volumes, hence including adversarial ones. A secret bit $b$ within the game determines whether $\Sigma_{0}$ (containing $N$ volumes) or $\Sigma_{1}$ (containing $N-1$ volumes) is first instantiated in the game and made to interact with the adversary. In both cases, we allow the adversary to choose the first $N-1$ passwords [^3] This represents the most unfavourable situation for the user, as we consider these passwords compromised anyway., and her goal is to guess $b$.
@@
    $\mathcal{A}$ chooses $N\in\left\{2,\ldots,N_{\mathsf{MAX}}\right\}$ and chooses $N-1$ passwords $P_{0},\ldots,P_{N-2}$.
@@
    A secret random bit $b\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\set{0,1}$ is drawn. If $b=0$, then an additional secret high entropy password is sampled $P_{N-1}\raisebox{-1.0pt}{$\,\xleftarrow{\raisebox{-1.0pt}{$\scriptscriptstyle\$$}}\,$}\set{0,1}^{\lambda}$, where $\lambda$ is the security parameter [^4] We abuse notation by representing a password as a binary string, but w.l.o.g. it is equivalent to the case of a user-chosen password with $\lambda$ bits of entropy. We assume that the user will choose a high entropy password at least for the hidden volume..
@@
    The game creates $N-b$ volumes: $\Pi.\mathtt{setup}(P_{0},\ldots,P_{N-1-b})\to\Sigma_{b}$
@@
      $\mathcal{A}$ chooses access patterns $O_{0}$ and $O_{1}$, where $O_{0}$ is, in the adversary’s intentions, aimed at $\Sigma_{0}$ (thus potentially containing some operations on $V_{N-1}$) and $O_{1}$ is aimed at $\Sigma_{1}$ (and so only contains operations on $V_{0},\ldots,V_{N-2}$).
@@
These constraints define the adversary goal, by specifying which two exact situations she has to distinguish between: if a PD scheme is secure (i.e., the adversary cannot distinguish) under the game enforcing such a constraint, the implication is that a user, having performed some access pattern $O_{0}$ including some operations on $V_{N-1}$, can plausibly claim to instead have executed a corresponding $O_{1}$, which only accesses the volumes $V_{0},\ldots,V_{N-2}$ (whose passwords have already been surrendered).
@@
Let us first clarify that some constraint is necessary in order to have any hope in the PD game against the adversary. Otherwise, the adversary could submit $O_{0}$ and $O_{1}$ containing completely different (logical) $\mathtt{write}$ accesses to the decoy volumes $V_{0},\ldots,V_{N-2}$, and there would be no way of making the two outcomes indistinguishable, since the adversary holds the passwords to those volumes, so it could trivially verify which of the two patterns was executed. This suggests the need for a minimal rule, stating that, whether $O_{0}$ or $O_{1}$ is executed, the resulting logical contents of the decoy volumes $V_{0},\ldots,V_{N-2}$ must be the same. From the user’s perspective, this basic requirement means that we do not try to disguise the $\mathtt{write}$s to decoy volumes as something else, both because we do not need to, and because there would be no way of doing it even if we wanted to.
Furthermore, we notice that many PD solutions (including those based on WoORAMs) treat $\mathtt{write}$ requests in a completely different way than $\mathtt{read}$ requests, i.e. a $\mathtt{write}$ request could trigger allocation or reshuffling of a certain volume sector. This might happen without breaking the minimal rule described above: if an adversary first $\mathtt{read}$s at a previously unallocated position $\beta_{\mathsf{L}}$, obtaining data $d$, and then $\mathtt{write}$s the same data $d$ at the same location $\beta_{\mathsf{L}}$, this might cause a change of state in the instance without changing its logical content (as mandated by the minimal rule above). This, in turn, would enable an attack where the adversary merely checks whether this state change happened or not. However, we must consider this attack trivial: detecting this kind of state change on a decoy volume would actually not compromise security. Therefore, in order to capture this concept, we also demand that the set of all blocks that are touched by $\mathtt{write}$ requests be the same for both $O_{0}$ and $O_{1}$, for all volumes $V_{0},\ldots V_{N-2}$. We stress that this extra constraint is also very minimal and does not weaken the security guarantees offered in practice; in fact, most schemes demand even stricter constraints, for example by demanding that $O_{0}$ and $O_{1}$ are of the same length, which we do not require here.